In September of 1995, McDonalds Corporation issued $150 million of senior notes due in 2005. The notes were issued at par and bore interest of 6 5/8%, paid semi-annually (i.e., interest of $33.125 per $1000, bond would be paid twice a year). The debt was rated AA by moodys. Interest payments on this debt were deductible for corporate tax purposes (you may assume that McDonald's marginal corporate tax rate was 35%), though principal repayments were not. All principal would be repaid in September 2005.

A. From McDonalds perspective, what is the effective after-tax cost of this debt (expressed as an annual percentage)?

B. How many dollars of taxes will mcdonalds save each year through the deduction of the interest expense on these notes from taxable income (you may assume that mcdonalds will have sufficient taxable income in future years to cover the interest expense of this debt)? What is the present value of these future tax savings?

Please provide a step by step approach i am completely lost on these problems.
Answer #1

A: Since bonds are issued at par, YTM of the bond = Coupon rate = 6.625%

After tax cost = 6.625%*(1-0.35) = 4.3063%

B: Amount of interest paid each semiannual period = Coupon*Total bond value/FV of bond

= 33.125*150000000/1000

= 4968750

Amount of tax saved each year = 35%*4968750*2 = 3478125

Present value of the amount assuming interest rate of 6.625% (Same as YTM)

PV of annuity = Annuity*(1-1/(1+rate)^number of terms)/rate

= 3478125*(1-1/(1+6.625%)^10)/6.625%

=24858038